Find the dimensions of the box described.

The length is 5 inches more than the width. The width is 2 inches more than the height. The volume is 264 cubic inches.

length ________in

width _________in

height ________in
Answer #1

let the length = x , width = y, height = z

The length is 5 inches more than the width

x=y+5

.

The width is 2 inches more than the height

y=z+2

and the volume is V=264 cubic inches.

.

volume is given by

V=xyz

V=\left(y+5\right)\left(z+2\right)z

264 =(2+2+5)(2+2)

264=\left(z+7\right)\left(z+2\right)z

264=z^3+9z^2+14z

z^3+9z^2+14z-264=0

here we have

a_0=264,\:\quad a_n=1

\mathrm{The\:dividers\:of\:}a_0:\quad 1,\:2,\:3,\:4,\:6,\:8,\:11,\:12,\:22,\:24,\:33,\:44,\:66,\:88,\:132,\:264

\mathrm{The\:dividers\:of\:}a_n:\quad 1

\mathrm{rational\:numbers:\quad \:}\pm \:\left(\frac{1,\:2,\:3,\:4,\:6,\:8,\:11,\:12,\:22,\:24,\:33,\:44,\:66,\:88,\:132,\:264}{1}\right)

\mathrm{rational\:numbers:\quad \:}\pm \:\left(1,\:2,\:3,\:4,\:6,\:8,\:11,\:12,\:22,\:24,\:33,\:44,\:66,\:88,\:132,\:264\right)

here z=4 is a root so take a synthetic division

z^2+13z+66=0

find discriminant

D=b^2-4ac

D=13^2-4\cdot \:1\cdot \:66

D=169-264

D=-95

here discriminant is negative so expression has complex roots

therefore our correct answer is

{\color{Red} z=4}

.

y=z+2

y=4+2

{\color{Red} y=6}

.

x=y+5

x=6+5

{\color{Red} x=11}

.

{\color{Red} length=11\:in,\:\:\:\:width=6\:in,\:\:\:\:\:height=4\:in}